# DG Singular equivalence and singular locus

### v5

@@ 3.2. Hochschild cohomology of DG singularity category
In this subsection, we give the proofs of the statements in the steps (1) and (2) in §1.3.
Our goal now is to endow $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ with a natural multiplication. First, by the construction of $F^{\ast}=\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$, $F^{\ast}$ has a natural multiplication structure from the endomorphism ring. Second, the following map
@@
induced by involution, gives a natural multiplication structure on $E^{\ast}$. Finally, there is a canonical $\Big{(}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),\mathrm{Bar}(R)\big{)}\Big{)}^{e}$-module structure on $\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. Therefore, viewing $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ as a semi-product $F^{\ast}\ltimes E^{\ast}$, we obtain on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ a natural multiplication structure.
Next, recall that the canonical morphism $E^{\ast}\rightarrow F^{\ast}$ of complexes is given by the following composition map:
@@
We next recall the support scheme of object in $D(R)$ and introduce the notion of the diagonal support scheme of $S_{sg}(R)$.
@@
Let $R$ be a commutative Noetherian ring. Let $R/I$ be the coordinate ring of singular locus on $\mathrm{Spec}(R)$. Then there is an algebra injection
$\iota_{I}:R/I\hookrightarrow\mathrm{HH}^{0}_{sg}(R).$
@@
First, there is a canonical DG algebra morphism
@@
which induces the following algebra homomorphism
@@
Next, the composition of algebra morphisms:
$R\xrightarrow{\pi_{R}}\mathrm{HH}_{sg}^{0}(R)\hookrightarrow\bigoplus_{i}\mathrm{HH}_{sg}^{i}(R)\cong\bigoplus_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$
gives the canonical $R$-module structure on $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$.
Since $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ contains the unit $\mathrm{id}$ and $\pi_{R}(1_{R})=\mathrm{id}$, its annihilator, as an $R$-module, is trivial. Hence, when we view $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ as a sheaf on $\mathrm{Spec}(R)$, to prove the proposition, it suffices to show that the support scheme of $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ is $V(I)\subseteq\mathrm{Spec}(R)$, which is equivalent to showing that the diagonal support scheme $\mathrm{DSupp}(S_{dg}(R))$ is exactly $V(I)$.
Let $\mathbf{p}$ be a point in $\mathrm{Spec}(R)$. We know that
@@
holds in $D(k)$, where $\mathrm{Bar}(R)\otimes_{R^{e}}R_{\mathbf{p}}^{e}$ is also a projective $R_{\mathbf{p}}^{e}$-module resolution of $R_{\mathbf{p}}$. On the other hand,
@@
in $D(k)$, where $\overline{\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$ is a flat resolution of $\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ over $R^{e}$. In the meantime, when we replace $R$ by $R_{\mathbf{p}}$, we obtain that
@@
in $D(R_{\mathbf{p}})$ as the argument in §2.1. Thus, we get
@@
Now we assume that $\mathbf{p}$ is a non-singular point. Since ${\mathbf{p}}$ is non-singular, $R_{\mathbf{p}}$ is a homologically smooth algebra. Then $\mathrm{Hom}_{S_{dg}(R_{\mathbf{p}}^{e})}(R_{\mathbf{p}},R_{\mathbf{p}})$ is trivial in $D(R_{\mathbf{p}})$. Hence $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{p}}$ is also trivial in $D(R_{\mathbf{p}})$. Thus, ${\mathbf{p}}$ is not contained in the diagonal support scheme $\mathrm{DSupp}(S_{dg}(R))$, which implies that
@@
Next, let $\mathbf{q}$ be a singular point in $\mathrm{Spec}(R)$, i.e., $\mathbf{q}\in V(I)$. As in above argument, we know that $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{q}}\cong\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$.
@@
Assume that $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is trivial in $D(R_{\mathbf{q}})$. It implies that $\mathrm{HH}^{0}_{sg}(R_{\mathbf{q}})$ is trivial algebra. From the distinguished triangle
$F^{\ast}_{\mathbf{q}}\rightarrow\mathrm{Cone}(E_{\mathbf{q}}^{\ast}\rightarrow F_{\mathbf{q}}^{*})\rightarrow E_{\mathbf{q}}^{\ast}[1],$
we have the long exact sequence
$\cdots\rightarrow\mathrm{H}^{0}(E_{\mathbf{q}}^{\ast})\rightarrow\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})\xrightarrow{\pi_{R}}\mathrm{H}^{0}\big{(}\mathrm{Cone}(E_{\mathbf{q}}^{\ast}\rightarrow F_{\mathbf{q}}^{*})\big{)}\rightarrow\mathrm{H}^{1}(E_{\mathbf{q}}^{\ast})\rightarrow\cdots,$
which induces a surjection
$\Theta:\mathrm{H}^{0}(E_{\mathbf{q}}^{\ast})\twoheadrightarrow\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})$
since
$\mathrm{H}^{0}\big{(}\mathrm{Cone}(E_{\mathbf{q}}^{\ast}\rightarrow F_{\mathbf{q}}^{*})\big{)}=\mathrm{HH}^{0}_{sg}(R_{\mathbf{q}})=0.$
Thus, for the identity element $\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}\in\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})$, we have
$\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}=\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}$
in $\mathrm{H}^{0}(F_{\mathbf{q}}^{\ast})$ for some $\alpha_{i}\in\left(\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),R_{\mathbf{q}}^{e})\right)^{r_{i}}=\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}},R_{\mathbf{q}}^{e})$ and $\beta_{i}\in\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}}$. Here, we use $(-)^{j}$ to denote the degree $j$ component of complex for any $j\in\mathbb{Z}$. For reader’s convenience, we view $\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}$ as an element in $F^{\ast}$. It follows that
$\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}-\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\in\mathrm{Im}(d_{F_{\mathbf{q}}^{\ast}}),$
where $d_{F_{\mathbf{q}}^{\ast}}$ is the differential of $F_{\mathbf{q}}^{\ast}$. Moreover, we have that
$\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\in\bigoplus_{1\leq i\leq m}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}\big{(}\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}},\mathrm{Bar}(R_{\mathbf{q}})^{-r_{i}}\big{)}.$
Let $l$ be an integer which is greater than all $r_{i}$. Let $N$ be an $R_{\mathbf{q}}^{e}$-module and $f\in\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}\mathrm{Bar}(R_{\mathbf{q}}),N[l]\big{)}$. Then $f$ is represented by a cycle, denoted by $\tilde{f}\in\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}})^{-l},N)$, in the complex $\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$. Thus we get the following composition
$\tilde{f}\circ\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}=0\in\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$
since $l\neq r_{i}$ for any $i$. In the meantime, we have that
$\tilde{f}\circ\Big{(}\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}-\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\Big{)}$
is in the image of the differential of $\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$ since
$\mathrm{id}_{\mathrm{Bar}(R_{\mathbf{q}})}-\Theta\big{(}\sum^{m}_{i=1}\alpha_{i}\otimes_{R_{\mathbf{q}}^{e}}\beta_{i}\big{)}\in\mathrm{Im}(d_{F_{\mathbf{q}}^{\ast}}).$
It implies that $\tilde{f}$ is in the image of the differential of $\mathbb{R}\mathrm{Hom}_{R_{\mathbf{q}}^{e}}(\mathrm{Bar}(R_{\mathbf{q}}),N)$. Thus, we obtain that $f=0$ in $\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}\mathrm{Bar}(R_{\mathbf{q}}),N[l]\big{)}$. It follows that
$\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}\mathrm{Bar}(R_{\mathbf{q}}),N[l]\big{)}\cong\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}\big{(}R_{\mathbf{q}},N[l]\big{)}=0.$
@@
The projective dimension of $R_{\mathbf{q}}$, as an $R^{e}_{\mathbf{q}}$-module, is less that $l$. It suggests that $R_{\mathbf{q}}$ is homologically smooth. Thus we obtain that $\mathbf{q}$ is a non-singular point in $\mathrm{Spec}(R)$. It is a contradiction to the fact that $\mathbf{q}$ is singular in $\mathrm{Spec}(R)$. Hence the assumption that
@@
Thus we obtain $V(I)=\mathrm{DSupp}(S_{dg}(R))$, that is, the diagonal support scheme of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ is $V(I)$. We thus completed the proof. ∎
Via the algebra morphism $\iota_{I}$, we view $\mathrm{HH}^{0}_{sg}(R)$ as an $R/I$-module.
@@
With the above setting, $\mathrm{HH}^{0}_{sg}(R)$ is a finitely generated $R/I$-module.
@@
First, we view $\mathrm{HH}^{0}_{sg}(R)$ as an $R$-module by the algebra morphism $\pi_{R}:R\rightarrow\mathrm{HH}^{0}_{sg}(R)$. Now, by Proposition 3.9, we have $\mathrm{HH}_{sg}^{0}(R)=\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}$.
Second, since $R$ is a Gorenstein Noetherian algebra, $R^{e}$ is also a Gorenstein Noetherian algebra. Hence, $R^{e}$ admits a bounded injective $R^{e}$-module resolution, denoted by $J^{*}$. There is a quasi-isomorphism of $R^{e}$-modules complexes:
@@
induced by the injective resolution $R^{e}\hookrightarrow J^{*}$ and another quasi-isomorphism of $R^{e}$-modules complexes:
@@
induced by Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$. Due to the boundedness of $J^{\ast}$, there are only finite many nontrivial cohomologies of $\mathbb{R}\mathrm{Hom}_{R^{e}}(R,J^{*})$.
In the meantime, for any $i\in\mathbb{Z}$, $\mathrm{Hom}_{K(R^{e})}(P^{*}_{R},R^{e}[i])\cong\mathrm{Hom}_{D(R^{e})}(R,R^{e}[i])$ as $R^{e}$-modules holds for any finitely generated projective $R^{e}$-module resolution $P^{*}_{R}$ of $R$, where $K(R^{e})$ is the homotopy category of complexes of $R^{e}$-modules. Hence we have that for any $i\in\mathbb{Z}$, $\mathrm{Hom}_{D(R^{e})}(R,R^{e}[i])$ is a finitely generated $R^{e}$-module.
By [12, Proposition 3.5], we get that $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$ is quasi-isomorphic to a bounded finitely generated $R^{e}$-module complex. Thus, $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$ admits a finitely generated projective $R^{e}$-module resolution, denoted by $P^{\ast}_{J^{*}}$. It suggests that there are quasi-isomorphisms
@@
It follows that
$R\otimes_{R^{e}}P^{\ast}_{J^{*}}\cong E^{*}$
@@
where $\mathrm{H}^{0}(F^{\ast})\cong\mathrm{HH}^{0}(R)\cong R$ and $\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}\cong\mathrm{HH}_{sg}^{0}(R)$. Since both $\mathrm{H}^{0}(F^{\ast})$ and $\mathrm{H}^{1}(E^{\ast})$ are finitely generated $R$-modules, we get that $\mathrm{HH}_{sg}^{0}(R)$ is also a finitely generated $R$-module from the long exact sequence.
Since $\pi_{R}$ factors through algebra morphism $\iota_{I}:R/I\hookrightarrow\mathrm{HH}_{sg}^{0}(R)$ (see Proposition 3.11), we obtain that $\mathrm{HH}_{sg}^{0}(R)$ is a finitely generated $R/I$-module. We thus completed the proof. ∎

### v6

@@ 3.2. Hochschild cohomology of DG singularity category
In this subsection, we give proofs of the statements in steps (1) and (2) in §1.3.
Our goal is to construct a multiplication on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$. First, we know that the composition of elements in the endomorphism ring $F^{\ast}=\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$ is a natural multiplication. Second, the following map
@@
induced by involution, gives a natural multiplication structure on $E^{\ast}$. Finally, there is a canonical $\Big{(}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),\mathrm{Bar}(R)\big{)}\Big{)}^{e}$-module structure on $\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. Therefore, viewing $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ as a semi-product $E^{\ast}\rtimes F^{\ast}$, we obtain a natural multiplication on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$.
Next, recall that the canonical morphism $E^{\ast}\rightarrow F^{\ast}$ is given by the following composition map:
@@
In the following, we recall the support scheme of object in $D(R)$ and the diagonal support scheme of $S_{sg}(R)$.
@@
Let $R$ be a commutative Noetherian ring. Let $R/I$ be the coordinate ring of the singular locus in $\mathrm{Spec}(R)$. Then there is an algebraic injection
$\iota_{I}:R/I\hookrightarrow\mathrm{HH}^{0}_{sg}(R)_{red},$
where $\mathrm{HH}^{0}_{sg}(R)_{red}$ is the quotient ring of $\mathrm{HH}^{0}_{sg}(R)$ by its nilpotent ideal.
@@
First, there is a canonical DG algebraic homomorphism
@@
which induces the following algebraic homomorphism
@@
Second, the composition of algebraic homomorphisms:
$\iota_{R}:R\xrightarrow{\pi_{R}}\mathrm{HH}_{sg}^{0}(R)\twoheadrightarrow\mathrm{HH}^{0}_{sg}(R)_{red}$
gives the canonical $R$-module structure on $\mathrm{HH}^{0}_{sg}(R)_{red}$.
Obviously $\iota_{R}$ induces an algebraic injection $R/I\hookrightarrow\mathrm{HH}^{0}_{sg}(R)_{red}$ when $\mathrm{Ker}(\iota_{R})\cong I$ as ideals of $R$. Now, we view $\mathrm{HH}^{0}_{sg}(R)_{red}$ as a sheaf on $\mathrm{Spec}(R)$. To prove the proposition, it suffices to show that the support scheme of $\mathrm{HH}^{0}_{sg}(R)_{red}$ is $V(I)\subseteq\mathrm{Spec}(R)$ since $V(\mathrm{Ker}(\iota_{R}))\cong\mathrm{Supp}\Big{(}\mathrm{HH}^{0}_{sg}(R)_{red}\Big{)}$.
Thus, it is enough to show that the diagonal support scheme $\mathrm{DSupp}(S_{dg}(R))$ is exactly $V(I)$. In fact, the support scheme $\mathrm{Supp}\Big{(}\mathrm{HH}^{0}_{sg}(R)_{red}\Big{)}$ is $\mathrm{Supp}\Big{(}\mathrm{HH}^{0}_{sg}(R)\Big{)}$. Moreover, $\mathrm{Supp}\Big{(}\mathrm{HH}^{0}_{sg}(R)\Big{)}$ is also $\mathrm{DSupp}(S_{dg}(R))$ by definitions, since the $R$-module structure on $\bigoplus_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ is given by the composition of $\pi_{R}$ and the natural injection $\mathrm{HH}^{0}_{sg}(R)\hookrightarrow\bigoplus_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$.
Let $\mathbf{p}$ be an element in $\mathrm{Spec}(R)$. We know
@@
in $D(k)$, where $\mathrm{Bar}(R)\otimes_{R^{e}}R_{\mathbf{p}}^{e}$ is a projective $R_{\mathbf{p}}^{e}$-module resolution of $R_{\mathbf{p}}$. On the other hand,
@@
in $D(k)$, where $\overline{\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$ is a flat resolution of $\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ over $R^{e}$. By the result, we obtain
@@
in $D(R_{\mathbf{p}})$, if $R$ is replaced by $R_{\mathbf{p}}$. Thus, we get
@@
Now assume $\mathbf{p}$ is a non-singular point, then $R_{\mathbf{p}}$ is a homologically smooth algebra, which implies that $\mathrm{Hom}_{S_{dg}(R_{\mathbf{p}}^{e})}(R_{\mathbf{p}},R_{\mathbf{p}})$ is trivial in $D(R_{\mathbf{p}})$. Hence $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{p}}$ is also trivial in $D(R_{\mathbf{p}})$. Thus, ${\mathbf{p}}$ is not contained in the diagonal support scheme $\mathrm{DSupp}(S_{dg}(R))$, which implies
@@
Next, let $\mathbf{q}$ be a singular point in $\mathrm{Spec}(R)$, i.e., $\mathbf{q}\in V(I)$. After similar arguments as above, we know $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{q}}\cong\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$.
@@
Assume that $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is trivial in $D(R_{\mathbf{q}})$. It implies that $\mathrm{HH}^{0}_{sg}(R_{\mathbf{q}})\cong\mathrm{Hom}_{D_{sg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is trivial algebra. Now, consider the surjective linear map:
$\mathrm{Hom}_{D_{sg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}}R_{\mathbf{q}})\otimes_{k}\mathrm{Hom}_{D_{sg}(R_{\mathbf{q}}^{e})}(F,R_{\mathbf{q}})\twoheadrightarrow\mathrm{Hom}_{D_{sg}(R_{\mathbf{q}}^{e})}(F,R_{\mathbf{q}}),$
which arises from the associativity of composition in the category $D_{sg}(R_{\mathbf{q}}^{e})$, for any object $F$ in $D_{sg}(R_{\mathbf{q}}^{e})$. Since $\mathrm{Hom}_{D_{sg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is is trivial, the left-hand side vanishes. Thus, we have
$\mathrm{Hom}_{D_{sg}(R_{\mathbf{q}}^{e})}(F,R_{\mathbf{q}})\cong 0$
for any object $F$ in $D_{sg}(R_{\mathbf{q}}^{e})$.
@@
By the Yoneda lemma, $R_{\mathbf{q}}\cong 0$ in $D_{sg}(R_{\mathbf{q}}^{e})$, which implies that $R_{\mathbf{q}}$ is homologically smooth. Thus we obtain that $\mathbf{q}$ is a non-singular point in $\mathrm{Spec}(R)$. It is a contradiction to the fact that $\mathbf{q}$ is singular in $\mathrm{Spec}(R)$. Hence the assumption that
@@
Then we obtain $V(I)=\mathrm{DSupp}(S_{dg}(R))$, which means that the diagonal support scheme of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ is $V(I)$. Thus, we complete the proof. ∎
Via the algebra homomorphism $\iota_{I}$, $\mathrm{HH}^{0}_{sg}(R)_{red}$ can be viewed as an $R/I$-module.
@@
With the above setting, $\mathrm{HH}^{0}_{sg}(R)_{red}$ is a finitely generated $R/I$-module.
@@
First, by Proposition 3.9, we have $\mathrm{HH}_{sg}^{0}(R)=\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}$.
Second, $R^{e}$ is also a Gorenstein Noetherian algebra, since $R$ is a Gorenstein Noetherian algebra. Hence, $R^{e}$ admits a bounded injective $R^{e}$-module resolution, denoted by $J^{*}$. There exist two quasi-isomorphisms
@@
via the injective resolution $R^{e}\hookrightarrow J^{*}$ and
@@
via the Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$. Due to the boundedness of $J^{\ast}$, there are only finite parts of the complex $\mathbb{R}\mathrm{Hom}_{R^{e}}(R,J^{*})$ have nontrivial cohomologies.
Meanwhile, $\forall i\in\mathbb{Z}$, $\mathrm{Hom}_{K(R^{e})}(P^{*}_{R},R^{e}[i])\cong\mathrm{Hom}_{D(R^{e})}(R,R^{e}[i])$ for any finitely generated projective resolution $P^{*}_{R}$ of $R$ as $R^{e}$-modules, where $K(R^{e})$ is the homotopy category of complexes of $R^{e}$-modules. Hence we have that for any $i\in\mathbb{Z}$, $\mathrm{Hom}_{D(R^{e})}(R,R^{e}[i])$ is a finitely generated $R^{e}$-module.
By [12, Proposition 3.5], we find that $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$ is quasi-isomorphic to a complex of finitely generated bounded $R^{e}$-modules. Thus, $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$ admits a finitely generated projective $R^{e}$-module resolution, denoted by $P^{\ast}_{J^{*}}$. It suggests that there are quasi-isomorphisms
@@
Thus, we have
$R\otimes_{R^{e}}P^{\ast}_{J^{*}}\xrightarrow{\sim}E^{*}$
@@
where $\mathrm{H}^{0}(F^{\ast})\cong\mathrm{HH}^{0}(R)\cong R$ and $\mathrm{H}^{0}\big{(}\mathrm{Cone}(E^{\ast}\rightarrow F^{*})\big{)}\cong\mathrm{HH}_{sg}^{0}(R)$. Since both $\mathrm{H}^{0}(F^{\ast})$ and $\mathrm{H}^{1}(E^{\ast})$ are finitely generated $R$-modules, we find that $\mathrm{HH}_{sg}^{0}(R)$ is also a finitely generated $R$-module from the long exact sequence. Thus, $\mathrm{HH}_{sg}^{0}(R)_{red}$ is a finitely generated $R$-module
Since $\iota_{R}$ factors through the algebraic homomorphism $\iota_{I}:R/I\hookrightarrow\mathrm{HH}_{sg}^{0}(R)_{red}$ (see Proposition 3.11), we obtain that $\mathrm{HH}_{sg}^{0}(R)_{red}$ is a finitely generated $R/I$-module. Eventually, we complete the proof. ∎